# Corpus Poisoning via Approximate Greedy Gradient Descent

### v1

@@ B.1 Candidate Set Quality
Table 4 shows the in-domain attack results of inserting 10 adversarial passages into the NQ and MS MARCO datasets, with a poison rate of around $0.0003\%$ of the NQ corpus and $0.0001\%$ of the MS MARCO corpus. Similar to injecting only 1 adversarial passage, we found that (1) The pretrained Contriever model is easy to attack: even with just 10 adversarial passages, all three baselines (AGGD, HotFlip and Random perturbation) successfully attack more than 90% of the queries in both NQ and MS MARCO datasets, tricking the Contriever into returning the adversarial passage among the top-1000 retrieved results. For NQ dataset, even random perturbation achieves a high ASR of 97.75%, while the other two methods using gradient information achieves higher ASR of $>99\%$. (2) AGGD still outperforms HotFlip on ANCE, with improvements of 11.95% and 25.58% on NQ and MS MARCO datasets, respectively. (3) Though attacking DPR-nq on NQ with only 1 adversarial passage does no better than random perturbation (as shown in Table 1), increasing the number of adversarial passage to 10 allows both AGGD and HotFlip to perform better than random perturbation, with ASR improvement of 10.08% and 6.67% over the Random perturbation baseline.
@@
NQ AGGD 99.46(0.01) 86.76(1.48) 85.53(0.42) 23.41(0.03) 16.87(0.07)
HotFlip 99.06(0.07) 86.72(2.25) 73.58(1.97) 25.77(1.32) 13.46(0.42)
Random 97.75(0.04) 78.06(2.56) 55.87(1.35) 15.34(0.74) 6.79(0.07)
MS MARCO AGGD 96.51(1.16) 54.65(1.16) 72.09(0.0) 13.95(0.0) 5.81(3.49)
HotFlip 96.51(3.49) 59.3(3.49) 46.51(2.33) 11.63(0.0) 2.33(0.0)
Random 91.86(1.16) 27.91(2.33) 30.23(9.3) 5.81(3.49) 1.16(1.16)
Table 4: In-domain attack success rate (ASR) of AGGD on NQ and MS MARCO dataset with 5 retrievers by injecting 10 adversarial passages. (adding 0.0003% corpus to NQ and 0.0001% to MS MARCO dataset). The best-performing attacking method is highlighted with in bold (higher ASR indicates better attack performance).
@@
FiQA-2018 NQ AGGD 90.97(0.08) 25.08(7.02) 12.27(0.39) 4.32(0.0) 2.7(0.85)
HotFlip 79.55(1.77) 20.14(2.86) 8.33(1.7) 1.77(1.31) 1.62(0.54)
Random 79.17(0.0) 14.43(4.55) 5.56(0.31) 1.23(0.0) 1.23(0.15)
MS MARCO AGGD 97.38(0.31) 45.99(25.31) 25.69(1.16) 21.22(7.95) 14.35(6.33)
HotFlip 94.06(1.47) 50.08(3.32) 18.75(0.39) 13.81(5.17) 20.52(0.77)
Random 83.41(0.23) 5.25(0.31) 8.33(0.93) 10.57(5.32) 4.71(3.01)
NFCorpus NQ AGGD 83.13(2.01) 43.5(2.32) 75.54(3.72) 25.23(6.97) 25.23(2.94)
HotFlip 88.85(0.62) 42.88(3.56) 63.0(1.39) 31.89(1.24) 27.09(0.16)
Random 68.27(0.46) 29.72(4.02) 48.3(2.48) 22.29(0.93) 21.21(2.01)
MS MARCO AGGD 97.06(0.15) 62.07(25.85) 95.2(0.46) 72.29(7.9) 64.24(4.49)
HotFlip 90.56(4.8) 78.02(9.6) 89.78(0.93) 69.35(11.15) 74.77(5.42)
Random 88.85(3.1) 63.78(1.86) 81.58(1.7) 47.21(3.87) 61.92(8.98)
Quora NQ AGGD 85.62(0.13) 35.84(2.36) 42.12(0.82) 2.7(0.32) 3.53(0.89)
HotFlip 77.47(1.67) 32.46(9.31) 33.01(2.55) 3.84(0.03) 4.04(0.13)
Random 67.66(5.7) 23.15(5.66) 26.36(0.42) 1.64(0.09) 2.36(0.47)
MS MARCO AGGD 91.8(2.98) 34.08(5.24) 48.4(0.52) 12.75(0.04) 12.0(1.43)
HotFlip 92.31(3.08) 40.83(3.2) 41.58(0.04) 12.93(0.48) 13.66(0.03)
Random 65.28(3.2) 11.79(3.52) 30.69(1.66) 6.43(0.13) 7.09(0.95)
SCIDOCS NQ AGGD 53.45(7.45) 45.7(17.7) 33.0(0.9) 1.75(0.35) 0.45(0.05)
HotFlip 49.7(2.2) 40.95(5.15) 25.4(5.7) 3.2(0.1) 1.0(0.2)
Random 37.7(4.1) 17.95(1.35) 13.25(0.95) 2.2(0.5) 0.55(0.25)
MS MARCO AGGD 65.4(2.5) 41.0(10.1) 45.25(2.15) 24.5(2.2) 6.4(0.0)
HotFlip 51.35(4.65) 67.55(8.25) 36.1(0.2) 21.85(2.95) 7.95(0.75)
Random 39.05(4.55) 13.6(4.1) 20.25(0.65) 14.7(2.9) 2.5(1.1)
SciFact NQ AGGD 66.83(5.5) 16.33(9.67) 56.83(9.5) 1.83(1.5) 0.33(0.33)
HotFlip 63.5(12.17) 15.67(3.0) 27.0(8.0) 2.83(1.17) 1.0(0.0)
Random 41.17(4.5) 5.0(3.0) 12.67(0.67) 1.0(0.0) 0.67(0.0)
MS MARCO AGGD 93.83(2.5) 49.83(13.17) 72.5(1.17) 17.33(5.67) 17.5(10.5)
HotFlip 89.5(3.17) 33.5(8.83) 58.83(6.17) 24.5(1.17) 21.17(10.83)
Random 80.83(8.17) 13.67(4.0) 32.67(8.0) 8.33(2.33) 14.5(2.17)
@@
Table 5: Out-of-domain top-50 attack success rate with 10 adversarial passages. Due to the computational constraints, results are averaged over 2 random runs, with standard deviations shown in parentheses. The combinations of attack and source dataset that achieve the highest ASR for each target domain and retriever are highlighted in bold.
@@ C.2 Retrievers
  • Dense Passage Retriever(DPR) Karpukhin et al. (2020) is a two-tower bi-encoder trained with a single BM25 hard negative and in-batch negatives. It has been used as the retrieval component of many Retrieval-Augmented Generation (RAG) models Lewis et al. (2020). In our paper, we use both the open-sourced Multi model (DPR-mul), which is a bert-base-uncased model trained on four QA datasets (NQ Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), WebQuestions Berant et al. (2013) and CuratedTREC Baudiš & Šedivỳ (2015)) and the single NQ model (DPR-nq).
@@ Appendix D Knowledge Poisoning Attacks to RAG
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims we made in the abstract accurately reflect the paper’s contributions and scope, and they are justified in the experiments section.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitation of the paper in the discussion section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not have theoretical result in this paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We disclose the details of the algorithm, as well as hyperparameters used to implement our algorithm.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We include the data and code in supplemental material to reproduce our results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specify the details in experiment setup section.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report confidence intervals for the experiments we made.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We list the compute resources required.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We confirm to the code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss broader impacts.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We don’t think our model has a high risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We properly credited and are the license and terms of use explicitly mentioned and properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We document our code.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We don’t do crowadsouring and research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We don’t do crowadsouring and research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@ B.1 Candidate Set Quality
Table 4 shows the in-domain attack results of inserting 10 adversarial passages into the NQ and MS MARCO datasets. Similar to injecting only 1 adversarial passage, we found that (1) The pretrained Contriever model is easy to attack: even with just 10 adversarial passages, all three baselines (AGGD, HotFlip and Random perturbation) successfully attack more than 80% of the queries in both NQ and MS MARCO datasets, tricking the Contriever into returning the adversarial passage among the top-20 retrieved results. For NQ dataset, even random perturbation achieves a high ASR of 80%, while the other two methods using gradient information achieves higher ASR of $>91\%$. (2) AGGD still outperforms HotFlip on ANCE, with improvements of 18.62% and 25.58% on NQ and MS MARCO datasets, respectively.
@@
NQ AGGD 91.57(3.5) 69.42(1.38) 82.08(2.01) 8.23(1.48) 1.9 (0.07)
Hot Flip 91.6(0.87) 57.37(5.95) 63.46(1.55) 5.08(0.1) 2.19(0.25)
Random 80.49(0.83) 33.73(2.27) 27.3(2.74) 3.94(1.22) 1.65 (0.06)
MS MARCO AGGD 83.72(4.65) 13.95(13.95) 93.02(0.0) 11.63(2.33) 5.81 (1.16)
Hot Flip 88.37(2.33) 13.95(11.63) 67.44(2.33) 10.46(1.16) 6.98(2.33)
Random 65.12(9.3) 1.16(1.16) 52.33(15.12) 8.14(1.16) 3.49 (1.16)
Table 4: In-domain attack success rate (ASR) of AGGD on NQ and MS MARCO dataset with 5 retrievers by injecting 10 adversarial passages). The best-performing attacking method is highlighted with in bold (higher ASR indicates better attack performance).
@@
FiQA-2018 NQ AGGD 92.67(0.23) 30.86(7.87) 19.52(1.47) 6.02(0.15) 4.71(1.47)
Hot Flip 84.1(0.77) 24.23(2.31) 13.35(2.7) 3.94(2.55) 3.24(1.23)
Random 83.1(0.39) 17.82(5.32) 8.26(0.54) 2.7(0.23) 2.31(0.46)
MS MARCO AGGD 97.69(0.31) 50.54(25.23) 37.19(0.31) 27.31(7.87) 20.6(7.64)
Hot Flip 95.22(1.08) 56.48(3.4) 28.16(2.08) 19.37(6.4) 28.47(1.31)
Random 86.73(0.15) 6.94(0.15) 13.27(1.39) 15.66(7.48) 7.25(4.63)
NFCorpus NQ AGGD 78.02(1.24) 35.76(3.25) 64.55(3.87) 18.11(4.8) 18.89(3.1)
Hot Flip 85.45(0.31) 37.0(3.87) 48.76(1.08) 20.43(0.31) 20.59(1.39)
Random 63.78(1.86) 24.46(3.41) 35.6(0.62) 16.72(1.55) 15.79(1.55)
MS MARCO AGGD 95.36(0.62) 57.9(26.01) 88.39(0.15) 57.89(8.05) 52.01(3.72)
Hot Flip 87.62(5.88) 70.28(12.69) 80.34(1.39) 59.75(10.84) 66.72(4.8)
Random 85.45(3.1) 51.39(5.57) 65.17(2.63) 34.98(1.24) 52.48(8.2)
Quora NQ AGGD 96.18(0.03) 63.14(3.93) 91.52(0.89) 20.96(3.63) 25.8(5.81)
Hot Flip 92.22(1.08) 57.88(10.45) 84.9(2.27) 24.82(0.15) 32.19(1.7)
Random 87.85(2.18) 46.28(7.64) 81.72(0.25) 15.84(0.97) 22.28(3.28)
MS MARCO AGGD 97.87(0.93) 61.72(7.29) 94.52(0.17) 45.92(1.0) 51.5(3.35)
Hot Flip 98.13(0.79) 69.88(3.63) 91.5(0.16) 43.86(2.64) 51.6(3.73)
Random 83.97(3.6) 35.24(4.43) 85.48(2.07) 32.88(0.31) 38.36(6.64)
SCIDOCS NQ AGGD 53.75(7.05) 46.4(17.7) 33.75(1.25) 1.7(0.3) 0.45(0.05)
Hot Flip 49.75(2.45) 40.7(5.4) 26.45(6.05) 3.35(0.15) 0.95(0.15)
Random 37.8(4.3) 18.25(1.75) 13.3(0.9) 2.3(0.4) 0.6(0.2)
MS MARCO AGGD 65.5(2.6) 41.65(9.85) 46.2(1.9) 24.9(2.2) 6.7(0.2)
Hot Flip 51.45(4.75) 68.1(8.3) 36.95(0.35) 22.4(2.9) 8.3(1.1)
Random 39.2(4.5) 13.9(4.3) 20.9(0.6) 15.55(3.35) 2.45(1.05)
SciFact NQ AGGD 60.33(6.0) 12.5(7.17) 39.33(9.33) 0.5(0.5) 0.17(0.17)
Hot Flip 58.83(12.83) 11.83(2.17) 16.5(3.5) 1.67(0.67) 0.5(0.17)
Random 34.33(4.33) 3.17(2.17) 8.17(1.5) 0.33(0.0) 0.5(0.17)
MS MARCO AGGD 91.5(4.5) 41.5(8.5) 58.17(0.83) 12.0(5.0) 12.0(7.67)
Hot Flip 85.83(5.17) 29.17(9.5) 42.33(4.0) 15.17(0.5) 14.67(7.33)
Random 75.5(9.5) 8.0(1.33) 21.5(6.5) 5.67(1.67) 10.33(1.0)
@@
Table 5: Out-of-domain top-20 attack success rate with 10 adversarial passages. Due to the computational constraints, results are averaged over 2 random runs, with standard deviations shown in parentheses. The combinations of attack and source dataset that achieve the highest ASR for each target domain and retriever are highlighted in bold.
B.2 In-domain attack success rate with different $k$.
Dataset Method Retriever
Contriever Contriever-MS ANCE DPR-mul DPR-nq
Top-1 NQ AGGD 85.06(5.06) 44.84(12.65) 16.36(1.52) 0.39(0.15) -
Hot Flip 83.77(1.43) 39.74(7.81) 8.66(0.62) 0.21(0.08) 0.03 (0.02)
Random 66.64(2.61) 8.82(0.87) 1.89(0.5) 0.07(0.04) 0.04(0.03)
MS MARCO AGGD 69.19(6.65) 11.05(12.67) 24.42(7.45) 1.16(1.16) 2.91(1.93)
Hot Flip 71.51(4.15) 12.79(8.47) 10.47(2.01) 2.33(1.64) 0.58 (1.01)
Random 52.33(6.26) 1.74(1.93) 5.23(4.47) - 0.58 (1.01)
Top-5 NQ AGGD 89.84(3.53) 55.63(10.96) 49.56(3.78) 1.9(0.55) 0.27 (0.16)
Hot Flip 88.69(0.72) 50.64(5.96) 31.36(1.16) 1.29(0.26) 0.28(0.1)
Random 75.28(1.8) 20.07(1.44) 10.12(1.82) 0.71(0.28) 0.24 (0.11)
MS MARCO AGGD 81.98(5.03) 20.35(16.72) 66.28(5.33) 8.72(3.02) 3.49(1.16)
Hot Flip 80.23(6.04) 19.77(10.2) 36.05(9.08) 3.49(1.16) 2.33 (0.0)
Random 62.21(7.6) 8.14(7.07) 20.93(4.03) 3.49(2.01) 0.58 (1.01)
Top-10 NQ AGGD 91.28(3.03) 59.39(9.83) 66.43(4.42) 3.62(1.02) 0.82(0.38)
Hot Flip 89.85(0.82) 54.52(5.41) 47.55(1.73) 2.69(0.24) 0.72 (0.2)
Random 77.8(1.45) 25.91(2.07) 18.8(3.26) 1.61(0.45) 0.66 (0.1)
MS MARCO AGGD 84.88(3.86) 22.67(17.35) 85.47(2.53) 9.3(3.68) 4.07(1.01)
Hot Flip 81.4(5.45) 20.35(10.58) 57.56(9.36) 4.65(1.64) 2.91 (1.01)
Random 65.12(7.89) 11.63(10.78) 35.47(8.28) 4.65(2.85) 1.16 (1.16)
Top-20 NQ AGGD 92.5(2.68) 63.45(8.68) 80.92(4.82) 6.88(1.72) 2.19(0.98)
Hot Flip 91.08(0.9) 58.43(4.53) 65.68(2.5) 5.4(0.36) 2.03 (0.25)
Random 80.24(0.92) 32.5(2.61) 31.0(4.3) 3.7(0.93) 1.66 (0.14)
MS MARCO AGGD 85.47(3.81) 24.42(17.44) 93.6(1.01) 12.79(3.86) 5.23 (1.01)
Hot Flip 83.72(5.93) 22.67(12.01) 76.16(9.21) 9.88(1.93) 5.81(2.01)
Random 66.86(7.24) 13.95(13.46) 49.42(11.2) 6.4(2.53) 2.91 (1.93)
Top-50 NQ AGGD 93.87(2.26) 68.69(6.66) 93.7(3.21) 15.59(3.41) 7.31(2.02)
Hot Flip 92.56(0.82) 63.28(3.64) 85.93(2.4) 13.37(1.1) 6.84 (1.1)
Random 83.01(0.46) 42.35(3.55) 53.01(5.31) 9.6(1.76) 5.46 (0.51)
MS MARCO AGGD 87.21(5.33) 31.98(20.3) 99.42(1.01) 22.67(5.55) 13.37(3.02)
Hot Flip 84.88(4.19) 32.56(18.16) 95.35(1.64) 19.77(3.49) 10.46 (5.07)
Random 69.19(4.47) 19.19(17.35) 75.58(5.33) 12.79(3.49) 4.65 (1.64)
Top-100 NQ AGGD 94.77(2.08) 72.59(5.28) 98.2(1.31) 27.94(5.7) 18.0(3.6)
Hot Flip 93.58(0.66) 67.27(2.97) 94.94(1.63) 24.48(2.26) 16.8 (1.79)
Random 85.11(0.41) 50.86(4.34) 71.57(5.81) 18.6(3.15) 12.81 (1.34)
MS MARCO AGGD 89.54(4.79) 40.12(24.36) 100.0(0.0) 41.28(5.03) 20.93(4.35)
Hot Flip 86.05(4.93) 40.12(18.12) 100.0(0.0) 31.98(5.04) 18.02 (5.78)
Random 70.93(2.6) 23.26(18.09) 90.12(6.01) 18.02(2.53) 9.88 (4.47)
Table 6: In-domain attack success rate (ASR) of AGGD on NQ and MS MARCO datasets with 5 retrievers by injecting 1 adversarial passage with varying $k=\{1,5,10,20,50,100\}$. Results are from 4 random runs with standard deviation in parenthesis.For the ease of presentation, we omit the results through ‘-’ if top-$k$ ASR is smaller than 0.1%.
B.3 Transferability over different retrievers.
Source Retriever Method Target Retriever
Contriever Contriever-MS ANCE DPR-mul DPR-nq
contriever AGGD 97.14 26.51 - - 1.2
Hot Flip 96.05 20.9 - 0.18 0.48
Random 90.5 11.8 - 0.14 1.39
contriever-msmarco AGGD 27.76 83.71 - - 0.28
Hot Flip 25.23 79.05 - 1.02 2.28
Random 26.24 78.38 - 2.05 7.16
ance AGGD 26.43 4.88 100.0 - -
Hot Flip 28.48 4.92 $9$
Random 9.86 6.39 99.3 - -
dpr-multi AGGD 15.78 3.67 - 87.84 22.47
Hot Flip 23.43 7.7 - 84.18 38.41
Random 18.43 4.15 - 79.35 10.41
dpr-single AGGD 10.85 9.99 - 29.16 89.46
Hot Flip 11.87 9.45 - 27.91 87.72
Random 22.02 7.37 - 20.5 82.55
Table 7: Attack transferability across models on NQ dataset.
@@ C.2 Retrievers
  • Dense Passage Retriever(DPR) Karpukhin et al. (2020) is a two-tower bi-encoder trained with a single BM25 hard negative and in-batch negatives. It has been used as the retrieval component of many Retrieval-Augmented Generation (RAG) models Lewis et al. (2020). In our paper, we use both the open-sourced Multi model (DPR-mul), which is a bert-base-uncased model trained on four QA datasets (NQ Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), WebQuestions Berant et al. (2013) and CuratedTREC Baudiš and Šedivỳ (2015)) and the single NQ model (DPR-nq).
@@ Appendix D Knowledge Poisoning Attacks to RAG
Appendix E Examples of the output passage from AGGD
Examples from attacking Contriever Examples from attacking contriever-MS
["##nius", "half", "melting", "definite", "[MASK]", "favor", "gma", "who", "should", "jew", "rain", "##pi", "##ntial", "upper", "kevin", "perrin", "##gren", "blew", "demonstrators", "remnants", "shelters", "that", "##ntial", "hue", "lest", "rainfall", "where", "rains", "rain", "heavy"] ["hostage", "tightly", "handful", "herbs", "where", "packets", "were", "##ssen", "dealers", "overnight", "symbol", "##atic", "##ised", "adventure", "bail", "##alis", "##kari", "[CLS]", "convicted", "ga", "##oit", "peace", "restore", "lifespan", "discrimination", "(", "maha", "##ter", "bank", "##ees"]
Table 10: Examples of adversarial passages